### v3

# Massless limit of massive self-interacting vector fields

Abstract
We study massive self-interacting vector field theories with mass added “by hand”. We show that the massless limit of the quartic self-interacting vector field theory is not smooth. Pathological behavior of the theory is not limited only at the classical level, even at the quantum level unitarity is violated at the two-loop. Using the Vainshtein mechanism, we show that it fails beyond the strong-coupling scale and hence a massless limit is not smooth even in quantum theory.
1 Introduction
The generation of mass in quantum field theories can be approached through two distinct methods: the Higgs mechanism Higgs1964 and the addition “by hand”. The Higgs mechanism Higgs1964, which is a cornerstone of the Standard Model, ensures renormalizability by spontaneously breaking gauge symmetries Lee1970; Boulware1970. In contrast, adding mass “by hand” often leads to non-renormalizable theories and can violate the Ward identities Ktorides.
Pathology at classical level: The Proca theory Proca1936, which describes a massive vector field, provides a well-understood framework for analyzing the possibility of a non-zero vector or gauge boson mass. The application of self-interacting vector field is also widely used in gravity and cosmology Esposito-Farese:2009wbc; DeFelice:2016yws; DeFelice:2016cri; Heisenberg:2017hwb; Kase:2017egk; Ramazanoglu:2017xbl; Annulli:2019fzq; Barton:2021wfj; Minamitsuji:2018kof; Herdeiro:2020jzx; Herdeiro:2021lwl, inflationary dynamics Dimopoulos:2006ms; Golovnev:2008cf; Maleknejad:2011jw; Maleknejad:2011sq; Adshead:2012kp; Emami:2016ldl; Garnica:2021fuu and dark matter theories Holdom:1985ag; Graham:2015rva; Agrawal:2018vin; Goodsell:2009xc; Bastero-Gil:2018uel; Ema:2019yrd; Kolb:2020fwh; Ahmed:2020fhc; Caputo:2021eaa; Antypas:2022asj. Problems in massive self-interacting vector fields where mass is added “by hand” are not only limited to the non-renormalizability of the theory. Recently a few groups have utilized hyperbolic formulation and have shown that interaction beyond the Proca limit, are not well-posed  [^1] A well-posed problem is characterized by the existence of a solution, uniqueness of the solution, and continuous dependence on initial conditions. due to the formation of singularities in 1+1 dimensions and the loss of hyperbolicity in higher dimensions PhysRevLett.129.151103; PhysRevD.108.044022. This loss of hyperbolicity implies that the classical equations of motion (EOM) change from hyperbolic (wave-like) to elliptical form, leading to ill-posedness PhysRevD.108.044022. We took the massless limit of these self-interacting theories and showed that they are also not smooth at the classical level.
The principle of continuity suggests that altering a physical theory should result in observables smoothly approaching their original values as the modification is reversed SchrodingerBass1955. However, massive self-interacting gauge theories with mass added “by hand” appear to violate this principle. It is fascinating how a minor modification in the theory drastically affects its behavior.
In this paper one of our primarily goal is to understand the fate of the pathology under the massless limit.
Quantum correction can avoid the classical singularity (kink) by smoothing it. The natural question is, does this pathological behavior persist at the quantum level? To answer this question we check unitarity of the theory in the massless limit at the scale where the equations of motion suffer from hyperbolicity loss.
Pathology at Quantum level: Conventional approaches indicate that these theories do not have a smooth massless limit, with perturbative series becoming singular in mass, suggesting a violation of unitarity or failure of the regular perturbation theory Hell2022. This issue is not unique to vector fields; linearized massive gravity also suffers from a similar problem known as the van Dam-Veltman-Zakharov (vDVZ) discontinuity vanDam1970. This discontinuity arises because the longitudinal mode in massive gravity fails to decouple from the other degrees of freedom as the mass approaches zero. Both General Relativity and massive gravity are nonlinear theories that are typically solved using perturbation theory. However, there exists a scale known as the Vainshtein radius Vainshtein1972gravity, at which nonlinear terms become significant, causing perturbation theory to break down for longitudinal modes Vainshtein1972; Vainshtein1972gravity. This breakdown leads to a strong coupling regime in which the longitudinal mode decouples, restoring General Relativity’s predictions in the massless limit Vainshtein1972gravity. Recent work attempted to demonstrate a smooth massless limit for massive self-interacting theories using the Vainshtein mechanism Hell2022. However, our findings using the hyperbolic formulation suggest that the massless limit is not smooth at the classical level, in presence of quartic self coupling. We revisited the non-covariant formulation and compared it with the hyperbolic formulation, which helps in determining the true behavior of the massless limit at the quantum level.
The plan of the paper is as follows: in section 2 we have introduced the non-linear Proca theory (NPL) with quartic self-interaction and discuss the hyperbolic formulation of its equation of motion. We show that the hyperbolicity of the equation is lost at a finite value of the field amplitude. Next in section LABEL:sec3 we discuss the perturbative classical study of the theory. In section LABEL:sec4, we present a non-covariant formulation of the theory, that is useful for the discussions in the later sections. Section LABEL:sec5 discusses the unitarity violation in the corresponding quantum theory and in sections LABEL:sec6 and LABEL:sec7 we describe the unitarity violation as an artifact of the breakdown of the corresponding perturbative approach. We explained the energy scales where the theory is well approximated. The implications of the massless limit has been discussed in all sections. The paper ends with a discussion in section LABEL:sec8. Necessary technicalities of the cubic and quartic NPL are presented in the first four appendices, whereas the last appendix contains results for massive Yang-Mills theory. Unlike massive abelian vector field theories, massive Yang-Mills, a theory of non-abelian vector field in which two types of intrinsic self-interactions are present: quartic and cubic derivative self interaction, has well posed classical EOM due to the internal symmetry. This has been recently shown in Ref. jose2025internalsymmetryrescuewellposed  [^2] This can be immediately verified by setting the additional interaction parameter to zero..
2 Hyperbolic Formulation of Nonlinear Proca Theory
The signature of the space-time metric is $(+,-,\dots,-)$.
In this section, we review the hyperbolic formulation for vector fields. We start with the Proca theory, which describes a massive vector field, and then proceed to self-interacting theories. We utilize the hyperbolic formulation to check the hyperbolicity of the equation of motion. The Lagrangian and the action for the Proca theory in a fixed space time with metric $g_{\mu\nu}$ is given by
${\cal{L}}_{PT}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+\frac{\mu^{2}}{2}X^{2}\ \quad\text{and,}\quad{\cal{S}}_{PT}=\int d^{4}x\sqrt{-g}\left(-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+\frac{\mu^{2}}{2}X^{2}\right)$ (1)
where $X^{2}=X_{\mu}X^{\mu}$ and $F_{\mu\nu}=\nabla_{\mu}X_{\nu}-\nabla_{\nu}X_{\mu}$ for the real vector field $X_{\mu}$. The corresponding field equation after utilizing the on-shell Lorenz condition ($\nabla_{\mu}X^{\mu}=0$) is
$g_{\mu\rho}\nabla^{\mu}\nabla^{\rho}X_{\nu}-R_{\mu\nu}X^{\mu}+\mu^{2}X_{\nu}=0;\quad\Box=g^{\mu\nu}\nabla_{\nu}\nabla_{\mu},$ (2)
where $R_{\mu\nu}$ is Ricci tensor. For Lorentzian space-time metric $g_{\mu\nu}$, the equation of motion (2) is hyperbolic (wave-like). In case of self-interacting theories, this is not true. Let us start with a massive vector field with quartic self-interaction
${\cal{L}}_{NPL}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}+\left(\frac{\mu^{2}}{2}X^{2}+\frac{\lambda}{4}\left(X^{2}\right)^{2}\right)\ ,$ (3)
where $X^{2}=X_{\mu}X^{\mu}$ and $F_{\mu\nu}=\nabla_{\mu}X_{\nu}-\nabla_{\nu}X_{\mu}$for the real vector field $X_{\mu}$. We call this Nonlinear Proca theory (NPL). The corresponding field equation after utilizing the on-shell Lorenz condition ($\nabla_{\mu}(zX^{\mu})=0$) is  PhysRevLett.129.151103
$\bar{g}_{\mu\rho}\nabla^{\mu}\nabla^{\rho}X_{\nu}$ $+2z^{\prime}z^{-1}\Theta_{\nu}-R_{\mu\nu}X^{\mu}-zX_{\nu}+\cdots=0$ (4)
$\text{where, }\bar{g}_{\mu\rho}$ $=g_{\mu\rho}+2z^{\prime}z^{-1}X_{\mu}X_{\rho},\Theta_{\nu}=X^{\mu}X^{\rho}\nabla_{\mu}F_{\nu\rho}$
$z$ $=-\left(\mu^{2}+\lambda X^{2}\right)\text{ and, }z^{\prime}=\frac{dz}{dX^{2}}.$ (5)
Where ellipses represent the product of single derivative terms. Since the EOM is a second-order differential equation, it can be classified based on the signature of the principal part of the differential equation. For the $1+1$ dimension, $\Theta_{\nu}=0$  PhysRevLett.129.151103, the EOM is controlled and classified solely by the effective metric $\bar{g}_{\mu\rho}$ and is given as,
$\bar{g}_{\alpha\beta}\nabla^{\alpha}\nabla^{\beta}X^{\mu}+\dots={\cal M}^{\mu}{}_{\alpha}X^{\alpha}$ (6)
where ${\cal M}_{\mu v}=-\mu^{2}\delta_{\mu v}+R_{\mu v}$ is the mass matrix. Let us now state two important definitions and a theorem that are necessary for our study.
Definition 1.
A Cauchy problem is called well-posed if there exist constants $K$ and $\alpha$ such that
[…]
Where $P(ik)$ is the principal symbol of a PDE Sarbach_2012. The significance of this definition is that it is based on the property that for each fixed time $t>0$ the norm $\left|e^{P(ik)t}\right|$ of the propagator is bounded by the constant $f(t):=Ke^{\alpha t}$, which is independent of the wave vector $k$ Sarbach_2012.